OpenStudy (anonymous):

A new photocopier can make 72 copies in 2 min. When an older photocopier is operational, the two photocopiers together can make 72 copies in 1.5 min. How long would it take the older photocopier to make 72 copies working alone?

OpenStudy (alekos):

just work out how many copies per minute for the new, N and h.m copies for old & new, O+N then subtract

OpenStudy (anonymous):

@Er.Mohd.AMIR

OpenStudy (anonymous):

i dont get it @alekos

OpenStudy (alekos):

N = 72/2 = 36 copies per minute O+N= 72/1.5 = ? copies per minute O=(O+N)-N = ? copies per minute

OpenStudy (alekos):

then to make 72 copies just divide this by the no. of copies/minute for O

OpenStudy (anonymous):

a little clearr?

OpenStudy (alekos):

I'll let you think about this for a few minutes